[Brahms-l] Shift report 20060618 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 18 2006 - 16:15:47 EDT
Shifter leader was hy
Shifters were sjs


1. We are at 90A700, and 2.3A1723. At the beginning of the shift, we had beam with low rate. After finishing the run from last shift, and a very short first run of out shift, we voted to dump the beam.  Around 10:30, we obtained a stable new store, but it did not last more than 1 hour, beam lost by unexpected quench link interlock. Beam came back at the very end of the shift, while ioserver and runserver were down, after fixing them, the data taking started. 

2. Runs achived: 
16219(from last shift), 
16220(short run before they dump the beam), 
16221, 16222(short one, beam was lost), 
16223(TPC pedestal run), 16224(Non-TPC pedestal run), 
16229 is going on on

3. None new detector problems so far

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 18 16:16:15 2006

This archive was generated by hypermail 2.1.8 : Sun Jun 18 2006 - 16:16:23 EDT